### v3

@@ 1 Introduction
[…]
@@ 2 Faster Two-Time-Scale Stochastic Gradient Method
[…]
@@
[…]
@@ 3.3 General Nonconvexity
$\min_{t\leq k}\mathbb{E}\left[\|\nabla h(\theta_{t})\|^{2}\right]$ $\leq\frac{32}{\alpha_{0}(k+1)^{1/2}}\left(\|\Delta f_{0}\|^{2}+\|\Delta g_{0}\|^{2}+(h(\theta_{0})-h^{\star})+y_{0}\right)+\frac{4\log(k+2)}{\alpha_{0}(k+1)^{1/2}}.$
@@ 4.1 Gradient-Based Policy Evaluation in Discounted-Reward MDP
$\nabla_{\theta}J(\theta)=-2\mathbb{E}_{\pi}[r(s,a)+\gamma\phi(s^{\prime})^{\top}\theta\phi(s)-\phi(s)^{\top}\theta\phi(s)]+2\gamma\mathbb{E}_{\pi}[\phi(s^{\prime})\phi(s)^{\top}]\,\omega=0,$ (20)
$\mathbb{E}_{\pi}[\phi(s)\phi(s)^{\top}]\,\,\omega=\mathbb{E}_{\pi}[r(s,a)+\gamma\phi(s^{\prime})^{\top}\theta\phi(s)-\phi(s)^{\top}\theta\phi(s)].$ (21)
It is straightforward to see that the objective (20)-(21) and stochastic oracles of the TDC algorithm exactly matches (1)-(3), where the operator in the lower level problem (21) is strongly monotone and the upper level objective is strongly convex. As a result, Algorithm 1 specializes to a variant of the TDC algorithm guaranteed to converge with rate $\widetilde{{\cal O}}(1/k)$, which matches the best known convergence rate of TDC under linear function approximation (Haque et al., 2023). When non-linear function approximation is considered, TDC solves a system of equations that resemble (20)-(21), where the lower level operator is still strongly monotone but the upper level objective is non-convex (Maei et al., 2009). The best known convergence rate for TDC under non-linear function approximation is $\widetilde{{\cal O}}(1/k^{-1/2})$ to a stationary point of $J$ (Wang et al., 2021), which we again match by our result in Theorem 3.
@@ 4.2 Policy Optimization for Linear-Quadratic Regulators
$\{u_{k}^{\star}\}\,\,=\,\,\min_{\{u_{k}\}}\,&\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}\Big{[}\sum_{k=0}^{T}\left(x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k}\right)\mid x_{0}\Big{]}\\ \text{subject to}\quad&x_{k+1}=Ax_{k}+Bu_{k}+w_{k},\end{aligned}$ (22)
@@
We are interested in finding the solution to (25) under unknown transition matrices $A$ and $B$, by taking a gradient-based approach. The natural gradient of $J$ with respect to $K$, which we denote by ~$\widetilde{\nabla}_{K}J$ and which can be regarded as $\nabla_{K}J$ on a slightly transformed system of coordinates, has the following closed-form expression
~$\widetilde{\nabla}_{K}J=2\left(R+B^{\top}P_{K}B\right)K-2B^{\top}P_{K}A.$
To obtain a reliable stochastic estimate of ~$\widetilde{\nabla}_{K}J$ the key is to estimate $R+B^{\top}P_{K}B$ and $B^{\top}P_{K}A$. For simplicity of notation, we define
@@
$\mathbb{E}_{x\sim\mu_{K},u\sim N(-Kx,\sigma^{2}I)}\left[M_{x,u,x^{\prime},u^{\prime}}\right]\left[\begin{array}[]{c}J(K)\\ \text{svec}(\Omega_{K})\end{array}\right]=\mathbb{E}_{x\sim\mu_{K},u\sim N(-Kx,\sigma^{2}I)}\left[c_{x,u}\right],$ (33)
@@
$M_{x,u,x^{\prime},u^{\prime}}=\left[\begin{array}[]{cc}1&0\\ \phi(x,u)&\phi(x,u)\left[\phi(x,u)-\phi\left(x^{\prime},u^{\prime}\right)\right]^{\top}\end{array}\right],\,\,c_{x,u}=\left[\begin{array}[]{c}x^{\top}Qx+u^{\top}Ru\\ \left(x^{\top}Qx+u^{\top}Ru\right)\phi(x,u)\end{array}\right].$
Equation (33) maps to the lower level problem (3). Under the assumption that $K$ is a stable controller with respect to $A$ and $B$ (Yang et al., 2019), the linear operator $\mathbb{E}[M_{x,u,x^{\prime},u^{\prime}}]$ is strongly monotone (Yang et al., 2019). The upper level problem that corresponds to (2) is to solve ~$\widetilde{\nabla}_{K}J=0$, which is known to observe the PL condition (Fazel et al., 2018). This means that we can apply Algorithm 1 to this problem, which specializes to an single-looped “actor-critic” algorithm with a finite-time convergence rate of $\widetilde{{\cal O}}(1/k)$ to the globally optimal solution measured by $J(K_{k})-J(\theta^{\star})$. The complexity of the current state-of-the-art single-looped algorithm for LQR is $\widetilde{{\cal O}}(1/k^{2/3})$, which we improve over. We also note the recent few works on nested-loop actor-critic algorithms that rely on batched multi-trajectory samples (Ju et al., 2022; Zhou and Lu, 2023). These algorithms achieve the complexity $\widetilde{{\cal O}}(1/k)$ as well, but in practice are much less convenient to implement than our algorithm which is single-looped and only requires a single online trajectory of samples.
@@
$g_{k+1}^{\Omega}$ $=(1-\lambda_{k})g_{k}^{\Omega}-\lambda_{k}\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]^{\top}\left(\left[x_{k}^{\top}\,\,\,u_{k}^{\top}\right]\hat{\Omega}_{k}\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]+\hat{J}_{k}-x_{k}^{\top}Qx_{k}-u_{k}^{\top}Ru_{k}\right)$
@@ A.6 Proof of Lemma 5
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
@@ B.1 Proof of Theorem 2
$\hskip 20.0pt+\mathbb{E}\Big{[}(1-\mu_{h}\alpha_{k})x_{k}-(\frac{\mu_{h}}{4}\alpha_{k}-\frac{6(L+1)^{5}}{\mu_{h}^{2}}\alpha_{k}^{2})x_{k}+(\frac{4L^{4}}{\mu_{h}^{3}}+2L)\alpha_{k}y_{k}+(\frac{2L^{2}\alpha_{k}}{\mu_{h}^{3}}+2L\alpha_{k}^{2})\|\Delta f_{k}\|^{2}$
@@
$\hskip 20.0pt-(\frac{\mu_{h}}{16}\alpha_{k}-\frac{30(L+1)^{6}}{\mu_{h}^{2}}\alpha_{k}\beta_{k})\mathbb{E}[x_{k}]-\left(\frac{\mu_{G}}{4}\beta_{k}-\left(14L^{3}+\frac{132(L+1)^{6}}{\mu_{h}^{3}}\right)\alpha_{k}-L^{2}\beta_{k}^{2}\right)\mathbb{E}[y_{k}],$
@@ B.5 Proof of Lemma 9
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right).$ (57)
@@ B.6 Proof of Lemma 10
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
@@ C.6 Proof of Lemma 15
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
@@
$\hskip 20.0pt+2L\alpha_{k}y_{k}+2L\alpha_{k}\|\Delta f_{k}\|^{2}+4L^{2}\alpha_{k}y_{k}+96L^{2}\alpha_{k}y_{k}+\frac{\alpha_{k}}{24}\|\nabla h(\theta_{k})\|^{2}+\frac{\mu_{G}\beta_{k}}{2}y_{k}+\frac{8\beta_{k}}{\mu_{G}}\|\Delta g_{k}\|^{2}$
@@ Appendix D Simulations Details
All algorithms start with zero initialization, i.e. $\theta_{0}=\omega_{0}=f_{0}=g_{0}=0$. The step sizes $\alpha_{k}$ and $\beta_{k}$ are set to constant values $45e-$ 4 and $32e-$ 3 for simplicity, while the step size $\lambda_{k}$ is selected as

### v4

@@ 1 Introduction
$\begin{aligned} \theta_{k+1}=\theta_{k}-\alpha_{k}F(\theta_{k},\omega_{k},X_{k}),\quad\omega_{k+1}=\omega_{k}-\beta_{k}G(\theta_{k},\omega_{k},X_{k}),\end{aligned}$ (7)
@@ 2 Faster Two-Time-Scale Stochastic Gradient Method
$\begin{aligned} \theta_{k+1}=\theta_{k}-\alpha_{k}f_{k},\quad\omega_{k+1}=\omega_{k}-\beta_{k}g_{k}.\end{aligned}$ (10)
@@
$\begin{aligned} f_{k+1}=(1-\lambda_{k})f_{k}+\lambda_{k}F(\theta_{k},\omega_{k},X_{k}),\quad g_{k+1}=\left(1-\lambda_{k}\right)g_{k}+\lambda_{k}G(\theta_{k},\omega_{k},X_{k}).\end{aligned}$ (11)
@@ 3.3 General Nonconvexity
$\min_{t\leq k}\mathbb{E}\left[\|\nabla h(\theta_{t})\|^{2}\right]$ $\leq\frac{32}{\alpha_{0}(k+1)^{1/2}}\left(\|\Delta f_{0}\|^{2}\hskip-1.0pt+\hskip-1.0pt\|\Delta g_{0}\|^{2}\hskip-1.0pt+\hskip-1.0pt(h(\theta_{0})\hskip-1.0pt-\hskip-1.0pth^{\star})\hskip-1.0pt+\hskip-1.0pty_{0}\right)+\frac{4\log(k+2)}{\alpha_{0}(k+1)^{1/2}}.$
@@ 4.1 Gradient-Based Policy Evaluation in Discounted-Reward MDP
$\nabla_{\theta}J(\theta)=-2\mathbb{E}_{\pi}\left[\Big(r(s,a)+\gamma\phi(s^{\prime})^{\top}\theta-\phi(s)^{\top}\theta\Big)\phi(s)\right]+2\gamma\mathbb{E}_{\pi}[\phi(s^{\prime})\phi(s)^{\top}]\,\omega=0,$ (20)
$\mathbb{E}_{\pi}[\phi(s)\phi(s)^{\top}]\,\,\omega=\mathbb{E}_{\pi}\left[\Big(r(s,a)+\gamma\phi(s^{\prime})^{\top}\theta-\phi(s)^{\top}\theta\Big)\phi(s)\right].$ (21)
It is straightforward to see that the objective (20)-(21) and stochastic oracles of the TDC algorithm exactly matches (1)-(3), where the operator in the lower level problem (21) is strongly monotone and the upper level objective is strongly convex. As a result, Algorithm 1 specializes to a variant of the TDC algorithm guaranteed to converge with rate $\widetilde{{\cal O}}(1/k)$, which matches the best known convergence rate of TDC under linear function approximation (Ma et al., 2020). When non-linear function approximation is considered, TDC solves a system of equations that resemble (20)-(21), where the lower level operator is still strongly monotone but the upper level objective is non-convex (Maei et al., 2009). The best known convergence rate for TDC under non-linear function approximation is $\widetilde{{\cal O}}(1/k^{-1/2})$ to a stationary point of $J$ (Wang et al., 2021), which we again match by our result in Theorem 3.
@@ 4.2 Policy Optimization for Linear-Quadratic Regulators
$\begin{aligned} \{u_{k}^{\star}\}\,\,=\,\,\min_{\{u_{k}\}}\,&\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}\Big[\sum_{k=0}^{T}\left(x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k}\right)\mid x_{0}\Big]\\ \text{subject to}\quad&x_{k+1}=Ax_{k}+Bu_{k}+w_{k},\end{aligned}$ (22)
@@
We are interested in finding the solution to (25) under unknown transition matrices $A$ and $B$, by taking a gradient-based approach. The natural gradient of $J$ with respect to $K$, which we denote by $\widetilde{\nabla}_{K}J$ and which can be regarded as $\nabla_{K}J$ on a slightly transformed system of coordinates, has the following closed-form expression
$\widetilde{\nabla}_{K}J=2\left(R+B^{\top}P_{K}B\right)K-2B^{\top}P_{K}A.$
To obtain a reliable stochastic estimate of $\widetilde{\nabla}_{K}J$ the key is to estimate $R+B^{\top}P_{K}B$ and $B^{\top}P_{K}A$. For simplicity of notation, we define
@@
$\hskip-8.5359pt\mathbb{E}_{x\sim\mu_{K},u\sim N(-Kx,\sigma^{2}I)}\left[M_{x,u,x^{\prime},u^{\prime}}\right]\left[\begin{array}[]{c}J(K)\\ \text{svec}(\Omega_{K})\end{array}\right]=\mathbb{E}_{x\sim\mu_{K},u\sim N(-Kx,\sigma^{2}I)}\left[c_{x,u}\right],$ (33)
@@
$M_{x,u,x^{\prime},u^{\prime}}\hskip-2.0pt=\hskip-2.0pt\left[\hskip-4.0pt\begin{array}[]{cc}1&0\\ \phi(x,u)&\hskip-4.0pt\phi(x,u)\left[\phi(x,u)-\phi\left(x^{\prime},u^{\prime}\right)\right]^{\top}\end{array}\hskip-6.0pt\right]\hskip-2.0pt,\,\,c_{x,u}\hskip-2.0pt=\hskip-2.0pt\left[\hskip-4.0pt\begin{array}[]{c}x^{\top}Qx+u^{\top}Ru\\ \left(x^{\top}Qx+u^{\top}Ru\right)\hskip-2.0pt\phi(x,u)\end{array}\hskip-7.0pt\right]\hskip-2.0pt.$
Equation (33) maps to the lower level problem (3). Under the assumption that $K$ is a stable controller with respect to $A$ and $B$ (Yang et al., 2019), the linear operator $\mathbb{E}[M_{x,u,x^{\prime},u^{\prime}}]$ is strongly monotone (Yang et al., 2019). The upper level problem that corresponds to (2) is to solve $\widetilde{\nabla}_{K}J=0$, which is known to observe the PL condition (Fazel et al., 2018). This means that we can apply Algorithm 1 to this problem, which specializes to an single-looped “actor-critic” algorithm with a finite-time convergence rate of $\widetilde{{\cal O}}(1/k)$ to the globally optimal solution measured by $J(K_{k})-J(\theta^{\star})$. The complexity of the current state-of-the-art single-looped algorithm for LQR is $\widetilde{{\cal O}}(1/k^{2/3})$, which we improve over. We also note the recent few works on nested-loop actor-critic algorithms that rely on batched multi-trajectory samples (Ju et al., 2022; Zhou and Lu, 2023). These algorithms achieve the complexity $\widetilde{{\cal O}}(1/k)$ as well, but in practice are much less convenient to implement than our algorithm which is single-looped and only requires a single online trajectory of samples.
@@
$g_{k+1}^{\Omega}$ $=(1-\lambda_{k})g_{k}^{\Omega}-\lambda_{k}\hskip-2.0pt\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]\hskip-3.0pt\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]^{\top}\hskip-7.0pt\left(\left[x_{k}^{\top}\,\,\,u_{k}^{\top}\right]\hat{\Omega}_{k}\hskip-2.0pt\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]\hskip-2.0pt+\hskip-2.0pt\hat{J}_{k}-x_{k}^{\top}Qx_{k}-u_{k}^{\top}Ru_{k}\right)$
@@ A.6 Proof of Lemma 5
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})\hskip-2.0pt-\hskip-2.0pt\omega^{\star}(\theta_{k})\|$
@@ B.1 Proof of Theorem 2
$\hskip 20.0pt+\mathbb{E}\Big[(1\hskip-2.0pt-\hskip-2.0pt\mu_{h}\alpha_{k})x_{k}\hskip-2.0pt-\hskip-2.0pt(\frac{\mu_{h}}{4}\alpha_{k}\hskip-2.0pt-\hskip-2.0pt\frac{6(L\hskip-2.0pt+\hskip-2.0pt1)^{5}}{\mu_{h}^{2}}\alpha_{k}^{2})x_{k}\hskip-2.0pt+\hskip-2.0pt(\frac{4L^{4}}{\mu_{h}^{3}}\hskip-2.0pt+\hskip-2.0pt2L)\alpha_{k}y_{k}\hskip-2.0pt+\hskip-2.0pt(\frac{2L^{2}\alpha_{k}}{\mu_{h}^{3}}\hskip-2.0pt+\hskip-2.0pt2L\alpha_{k}^{2})\|\Delta f_{k}\|^{2}$
@@
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]\hskip-2.0pt+\hskip-2.0pt(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]+(1-\mu_{h}\alpha_{k})\mathbb{E}[x_{k}]\hskip-2.0pt+\hskip-2.0pt(1-\mu_{G}\beta_{k})\mathbb{E}[y_{k}]\hskip-2.0pt+\hskip-2.0pt4B\lambda_{k}^{2}$
@@
$\hskip 20.0pt-(\frac{\mu_{h}}{16}\alpha_{k}-\frac{30(L\hskip-2.0pt+\hskip-2.0pt1)^{6}}{\mu_{h}^{2}}\alpha_{k}\beta_{k})\mathbb{E}[x_{k}]-\left(\frac{\mu_{G}}{4}\beta_{k}\hskip-2.0pt-\hskip-2.0pt\left(14L^{3}\hskip-2.0pt+\hskip-2.0pt\frac{132(L\hskip-2.0pt+\hskip-2.0pt1)^{6}}{\mu_{h}^{3}}\right)\alpha_{k}\hskip-2.0pt-\hskip-2.0ptL^{2}\beta_{k}^{2}\right)\mathbb{E}[y_{k}],$
@@ B.5 Proof of Lemma 9
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}\hskip-2.0pt+\hskip-2.0ptLy_{k}\hskip-2.0pt+\hskip-2.0pt\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right).$ (57)
@@ B.6 Proof of Lemma 10
$\leq 2\|\omega_{k}\hskip-2.0pt-\hskip-2.0pt\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})\hskip-2.0pt-\hskip-2.0pt\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
@@ C.6 Proof of Lemma 15
$\leq 2\|\omega_{k}\hskip-2.0pt-\hskip-2.0pt\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
@@
$\hskip 20.0pt+\hskip-2.0pt2L\alpha_{k}y_{k}+2L\alpha_{k}\|\Delta f_{k}\|^{2}\hskip-2.0pt+\hskip-2.0pt4L^{2}\alpha_{k}y_{k}+96L^{2}\alpha_{k}y_{k}\hskip-2.0pt+\hskip-2.0pt\frac{\alpha_{k}}{24}\|\nabla h(\theta_{k})\|^{2}\hskip-2.0pt+\hskip-2.0pt\frac{\mu_{G}\beta_{k}}{2}y_{k}\hskip-2.0pt+\hskip-2.0pt\frac{8\beta_{k}}{\mu_{G}}\|\Delta g_{k}\|^{2}$
@@ Appendix D Simulations Details
All algorithms start with zero initialization, i.e. $\theta_{0}=\omega_{0}=f_{0}=g_{0}=0$. The step sizes $\alpha_{k}$ and $\beta_{k}$ are set to constant values 5e45e-4 and 2e32e-3 for simplicity, while the step size $\lambda_{k}$ is selected as